food goblin supermarkets use both cashiers and baggers to serve customers at check out. during the first 5 hours of each workday (monday–friday), 4 cashiers and 3 baggers serve approximately 17 customers per hour. a cashier and a bagger who require approximately 4 minutes at checkout and 2 minutes at bagging serve each customer.

a. the utilization of cashiers is 28.5%
The utilization of baggers is 18,7%

B. Assume that both baggers and cashiers are cross- trained to perform both activities so that they can serve customers independently. Customers are now both checked and their groceries bagged by one individual. Further, assume that it takes 8 minutes for one individual to both cash cash out and bag each customer’s groceries. Calculate the utilization of this new group of 7 cross- trained employees.
The utilization of a cross-trained employees is ( )% ( Enter your response rounded to one decimal place.)
Answer #1

a has been already solved hence doing the calculation for b)

b)

Output - 17 customers per hour

Total Number of Manpower - 7

Unit Customer Service Time - 8 minutes

Per Hour Number of Customer Served = 60/8

Total Customers Served by 7 employees = 7*(60/8)

Utilization = Output/Total Customer Served by 7 employees

Utilization = 17/(7*(60/8))

Utilization = 32.38%

Utilization = 32.4%